# No Vulnerability Data, No Problem: Towards Predicting Mean Time To Remediate In Open Source Software Dependencies

### v2

@@ Abstract.
Timely remediation of vulnerabilities in software dependencies is critical for the security of the software supply chain. As such, researchers have proposed tools and metrics to help practitioners assess the security practices of each of their dependencies. Conceptually, a dependency-focused Mean-Time-To-Remediate (MTTR) metric can provide a historical perspective on how long it takes a given package to update vulnerable versions of its dependencies. However, existing MTTR metrics focus on a package fixing bugs in its own code, not its dependencies. Simultaneously, existing dependency update metrics do not aggregate values for the entire package and are not sensitive to aspects important for vulnerabilities (e.g., floating version constraints). The goal of this study is to aid industry practitioners, including developers, in assessing the risk of dependencies through a novel metric approximating mean-time-to-remediate vulnerabilities in their dependencies that is evaluated by an empirical study. We propose a novel algorithm for computing MTTR called MTTRdep and a companion metric called Mean-Time-To-Updatedep (MTTUdep), which considers all version updates, including vulnerability fix updates. We conduct a large-scale study using $163,207$ packages in npm, PyPI, and Cargo, of which only $22,513$ packages produce MTTRdep because of the lack of vulnerability data. We further study how package characteristics (e.g., contributors and version counts) influence MTTUdep and MTTRdep and explore how long packages retain outdated vulnerable dependencies in npm, PyPI, and Cargo. Our results indicate that industry practitioners can reliably use MTTUdep as a proxy for MTTRdep when available vulnerability data is insufficient.
@@ 1. Introduction
Vulnerable dependencies are widely present in both open-source software (OSS) and proprietary codebases. According to the Synopsys 2025 “Open Source Security and Risk Analysis Report” (blackduck, ), 86% of codebases contain at least one vulnerable open source dependency, and 81% of codebases contain high or critical risk vulnerabilities resulting from dependencies. The majority of codebases had vulnerable dependencies for more than two years despite the availability of a fixed version (ossf-scorecard, ; blackduck-2021, ). This delay occurs because of fear of breaking changes and the cost associated with updating vulnerable dependencies to fixed versions (derrKeepMeUpdated2017, ).
@@
Industry practitioners desire metrics that help assess the trustworthiness, risk, and security practices of open-source dependencies. For example, OpenSSF Scorecard (ossf-scorecard, ) considers 18 security practices to assess a package, including the “maintained” check, which determines if the project is maintained by checking activity in the last 90 days, and the “vulnerabilities” check detects if there are unfixed known vulnerabilities in the project or its dependencies.
@@
Mean-Time-To-Remediate (MTTR) is also often discussed in the context of vulnerability management (sonatype-2019, ). In contrast to Scorecard’s ‘maintained’ metrics, MTTR provides a historical perspective. However, MTTR has largely been discussed in the context of a project fixing its own vulnerabilities, not vulnerabilities in dependencies (gokhaleTimeStructureBased1999, ; calvoApplyingGoalQuestion2023, ; johnsonSurveySoftwareTools1988, ; morrisonMappingFieldSoftware2018, ). Dependency update metrics such as technical lag (gonzalez-barahona_technical_2017, ) consider management of dependency versions; however, they have not been applied to vulnerabilities. Conceptually, a lower value in a dependency update metric suggests faster and more consistent updates, signaling ongoing maintenance and reliability (sonatype-2024, ), which could help assess the security risks of adopting a package. However, the connection between dependency updates and exposure to vulnerabilities in dependencies has not yet been demonstrated.
Practitioners assessing the risk associated with using a given dependency face two key challenges when considering how it manages vulnerabilities in its dependencies.
1. Insufficient Vulnerability Data. Currently, only 2.4K of npm packages have known vulnerabilities, while the ecosystem hosts over 3.27M packages. Practitioners can only compute MTTR for the dependents of those 2.4k packages.
@@
2. Metric Limitations. The calculations used by existing dependency update metrics often do not handle floating version constraints or are not designed to handle vulnerabilities (coxMeasuringDependencyFreshness2015, ; gonzalez-barahona_technical_2017, ). Existing metrics also often calculate an update time for each of a dependency and for each version of a package, making it difficult to understand that package’s update practices as a whole. Finally, there are no existing publicly available dependency update metrics specifically for vulnerabilities. While some industry reports use MTTR in this way, their specific calculations are not available.
The goal of this study is to aid industry practitioners, including developers, in assessing the risk of dependencies through a novel metric approximating mean-time-to-remediate vulnerabilities in their dependencies that is evaluated by an empirical study. Specifically, we hypothesize that a package’s dependency update metric (Mean-Time-To-Update, MTTU) can serve as an indicator of how quickly the package would remediate security vulnerabilities (MTTR) in dependencies, even if the package never had a vulnerable dependency before. We also overcome challenges with dependency update calculations and propose novel algorithms for computing MTTU and MTTR, which we denote MTTUdep and MTTRdep, respectively.
We use MTTUdep and MTTRdep to answer the following research questions: RQ1:Can MTTUdep be used as a proxy for MTTRdep? RQ2:How do package characteristics influence MTTUdep and MTTRdep? RQ3:How long do packages have outdated dependencies, and how long do packages have vulnerable dependencies?
To answer these questions, we first collect package version release information and dependency relations of packages from the npm, PyPI, and Cargo ecosystems and compute MTTUdep and MTTRdep for $163,207$ packages. For RQ1, we perform an equivalence analysis using TOST (two one-side t-tests) and analyze correlations between MTTUdep and MTTRdep. For RQ1, we use correlation tests to understand the association between various package characteristics (e.g., contributors count, version count). We further conduct regression analysis to quantify which package characteristics matter more in influencing MTTUdep and MTTRdep values. Finally, for RQ1, we analyze the distributions of MTTUdep and MTTRdep to explore how long packages retain outdated and vulnerable dependencies in the ecosystems.
@@
Contributions. In summary, this paper contributes (1) a detailed algorithm and process for quantifying the dependency update practice of a package using our novel metrics; (2) statistical hypothesis testing on using MTTUdep as a proxy for MTTRdep, when vulnerability data is not available; (3) a large-scale analysis of the dependency update metrics in npm, PyPI, and Cargo packages; (4) correlation and regression analysis illustrating which package characteristics impact the likelihood of higher MTTUdep and MTTRdep; and (5) actionable insights for open-source developers and supply chain security researchers based on our findings.
@@
We provide our replication package in Zenodo (zenodo-artifact, ) (available for reviewers till Oct’25). Upon acceptance of the paper, we will make it public.
2. Related Work
@@
In this section, we provide a brief overview of the related work.
@@
Further research by Zerouali et al. (zeroualiFormalFrameworkMeasuring2019, ) propose a formal framework for measuring technical lag in software ecosystems. They analyzed 4M releases of 500K npm packages considering the evolution of technical lag over time. They found that technical lag induced by direct dependencies in npm packages increases over time due to missed updates, including major releases. Stringer et al. (stringerTechnicalLagDependencies2020, ) study the technical lag of dependencies in a large-scale cross-ecosystem fashion containing packages from 14 package managers. They found that pinned dependencies are the main reason behind technical lag. Zerouali et al. (zeroualiMultidimensionalAnalysisTechnical2021, ) expand the idea of ‘technical lag’ into multiple dimensions (package lag, time lag, version lag, vulnerability lag, and bug lag) and study the technical lag in 140K Docker images. They found that the median time lag of community Docker images is over a year. Although previous studies have explored technical lag in terms of general dependency updates, understanding vulnerable dependency update practice was not their goal. In contrast, we study the dependency update metric and vulnerable dependency update metric and investigate their relationship.
@@
Pashchenko et al. (pashchenkoVulnerableOpenSource2018, ) studied the most used Java dependencies in SAP software and found that only updating the dependencies’ version can remove 81% vulnerable dependencies. Kula et al. (kulaDevelopersUpdateTheir2018, ) studied the update behavior of developers w.r.t. security advisories and found that developers do not update their vulnerable dependencies regularly. Kumar et al. (kumar_comprehensive_2024, ) conducted a study to understand how widespread vulnerabilities are and how quickly they are being fixed. They found that for most programming languages, a critical vulnerability persists on average for over a year before being fixed. Studies in outdated dependencies and vulnerable dependencies are focused on either all updates or only security updates but not both. In contrast, we focus on both outdated and vulnerable dependencies and explore their relationship using our proposed metrics.
MTTU and MTTR. The metrics MTTU and MTTR have been used in the software reliability and maintenance domain for a long time (gokhaleTimeStructureBased1999, ; calvoApplyingGoalQuestion2023, ; johnsonSurveySoftwareTools1988, ; morrisonMappingFieldSoftware2018, ). Researchers have also studied different security metrics (time to close bug/vulnerability, window of exposure, vulnerability count) (davidsonGoodBadUgly2009, ; beresUsingSecurityMetrics2009, ) of various categories (time metric, vulnerability metric) in the software security domain, but these metrics are focused on packages having the vulnerability. In our study, we focus on measuring dependency update metrics for packages having vulnerable dependencies, not the vulnerable package itself.
@@ 3. Challenges Applying Existing Update Metrics
$0.9.2:\operatorname{2022-03-22}$ ${}^{\wedge}3.4.0$ $3.4.0:\operatorname{2022-03-18}$ $3.4.0:\operatorname{2022-03-18}$ 00
@@ 3.1. Technical Lag Using tLag
Time lag ($tLag$(zerouali_empirical_2018, ; decan_evolution_2018, ) is a way to measure technical lag to assess the outdatedness of a dependency in terms of time. Conceptually, $tLag$ measures the time difference between the release time of the latest version of a dependency and the release time of the version of the dependency used by a package, at the time the package was released.
Let $pkg$ be a package, and $pkg_{v}$ indicate specific version $v$ of $pkg$. Let $dep$ be a direct dependency of $pkg$. When specifying $dep$ as a dependency, $pkg_{v}$ specifies a dependency constraint $pkg_{v}.dep.\bar{c}$. Package managers use dependency constraints to resolve the highest version that satisfies the constraint. For example, a dependency constraint $pkg_{v}.dep.\bar{c}$ of “${}^{\wedge}1.2.3$” is satisfied for the range [¿=1.2.3, ¡ 2.0.0]. We use the function $lastAllowed(dep,\bar{c},t)$ to denote the version resolution of $dep$ for constraint $\bar{c}$ at time $t$. We further use the function $latest(dep,t)$ to denote the latest version of $dep$ at time $t$ and $time(\cdot)$ to denote the time a given version is released. Hence, the $tLag$ of $pkg_{v}$ for $dep$ is formally defined as follows:
@@
Table 1 shows an example computation of technical lag for codemod-cli’s dependency on simple-git. The first two columns of Table 1 indicate the version with the date of different releases of package codemod-cli and the version constraint specified for dependency simple-git. The “lastAllowed(simple-git)” column, represents the resolved version of dependency simple-git with specified constraint at the time of the release of codemod-cli. For computing the technical lag, we need to know the latest available version of the dependency simple-git at the time of the release of each version of codemod-cli according to Eq. 1, which represents in column “latest(simple-git)”. The tLag column is the computed technical lag. We subtract the release date of lastAllowed(simple-git) from the release date of latest(simple-git) to obtain the technical lag for each release of codemod-cli. The computed tLag of five versions of codemod-cli were 707, 78, 78, 100, and 0, consecutively, in Table 1 and Fig. 1.
@@ 3.2. Design Gaps
Package-Level Metric. A dependency update metric should provide a way to aggregate all data points into one data for each package so that developers can make sense of the metrics and use them for comparison if desired. One of the findings of Wermke et al. (wermkeAlwaysContributeBack2023, ) was that developers’ dependency selection metrics and criteria focus on quickly accessible numbers and facts, such as downloads, GitHub stars, and time since last release, substantiating the necessity of having easily comprehensible metrics. However, tLag computes the technical lag for each version of the package for a dependency as shown in Table 1. For developers, understanding what to do with all of these tLag values for a package with multiple releases and multiple dependencies is hard. Having an aggregation would improve the understandability of dependency update metrics.
@@
Handling Vulnerabilities. A security-oriented dependency update metric should incorporate the update characteristics of vulnerable dependencies. Technically, tLag’s measurement can be modified to only consider the dependency’s security fixes adopted by the package in its formal framework. However, no previous work measured tLag for measuring vulnerable dependency update practices.
4. Proposed Dependency Update Metrics

### v3

# How Quickly Do Development Teams Update Their Vulnerable Dependencies?

@@ Abstract.
Industry practitioners are increasingly concerned with software that contains vulnerable versions of third-party dependencies that are included both directly and transitively. To address this problem, projects are encouraged to both (a) quickly update to non-vulnerable versions of dependencies and (b) be mindful of the update practices of the dependencies they choose to use. To this end, researchers have proposed metrics to measure the responsiveness of the development teams of the packages in keeping their dependencies updated: Mean-Time-To-Update (MTTU) and Mean-Time-To-Remediate (MTTR). While MTTU covers all dependencies, MTTR quantifies the time needed for a package to update its vulnerable dependencies. However, existing metrics fail to capture important nuances, such as considering floating versions and prioritizing recent updates, leading to inaccurate reflections of a development team’s update practices. The goal of this study is to aid practitioners in understanding how quickly packages update their dependencies. We propose two novel metrics, Mean-Time-To-Update for dependencies (MTTUdep) and Mean-Time-To-Remediate for vulnerable dependencies (MTTRdep), that overcome the limitations of existing metrics. We conduct an empirical study using $163,207$ packages in npm ($117,129$), PyPI ($42,777$), and Cargo ($3,301$) and characterize how the ecosystems differ in MTTUdep and MTTRdep, as well as what package characteristics influence MTTUdep and MTTRdep. We found that most packages have a relatively fast dependency update practice. We also found that older packages tend to have higher MTTUdep and MTTRdep values. We further study whether MTTUdep can be used as a proxy for MTTRdep when sufficient vulnerability data is not available. As we did not find enough statistical evidence for a strong proxy, our findings suggest that MTTUdep could only be partially used (may be used but with caution) as a proxy for MTTRdep when vulnerability data is not available. This latter finding is particularly important given that only 1363 npm (0.04%), 694 PyPI (0.11%), and 383 Cargo (0.20%) packages have reported vulnerabilities, and the existence of MTTUdep will allow practitioners to make more informed decisions about the dependencies they choose.
@@ 1. Introduction
Vulnerable dependencies are widely present in both open-source software (OSS) and proprietary codebases. According to the Synopsys 2025 “Open Source Security and Risk Analysis Report” (blackduck, ), 86% of codebases contain at least one vulnerable open source dependency, and 81% of codebases contain high or critical risk externally reported vulnerabilities resulting from dependencies. The majority of codebases had vulnerable dependencies for more than two years despite the availability of a fixed version (ossf-scorecard, ; blackduck-2021, ). This delay occurs because of fear of breaking changes and the cost associated with updating vulnerable dependencies to fixed versions (derrKeepMeUpdated2017, ).
@@
This issue of vulnerable dependencies highlights the industry practitioners’ need for metrics to measure the responsiveness of development teams to updating their open-source dependencies. For example, the OpenSSF Scorecard (ossf-scorecard, ) evaluates a package based on 18 security practices. Among these practices are the “maintained” check, which determines if the project is maintained by checking activity in the last 90 days, and the “vulnerabilities” check, which detects if there are unfixed externally reported vulnerabilities in the project or its dependencies.
@@
While Scorecard’s “maintained” metric focuses on recent activity, other metrics, such as Mean-Time-To-Update(MTTU) and Mean-Time-To-Remediate (MTTR), provide a historical perspective on how long it takes for the development team of a package to update their dependencies. MTTU captures the time to update all dependency versions, while MTTR focuses specifically on the time to update vulnerable dependency versions to the fixed version (sonatype-2019, ). MTTR has been extensively discussed in the context of a project fixing its own vulnerabilities, but not vulnerabilities in dependencies (gokhaleTimeStructureBased1999, ; calvoApplyingGoalQuestion2023, ; johnsonSurveySoftwareTools1988, ; morrisonMappingFieldSoftware2018, ). Dependency update metrics, such as technical lag (gonzalez-barahona_technical_2017, ), consider management of dependency versions; however, they have not been applied to vulnerabilities. Furthermore, active maintenance (e.g., regular dependency update) is a desired criterion for developers in selecting a dependency as highlighted by Vargas et al. (larios_vargas_selecting_2020, ). In this context, a lower value in a dependency update metric suggests faster and more consistent updates, signaling ongoing maintenance and reliability (sonatype-2024, ).
Practitioners measuring the responsiveness of a team in updating its vulnerable and outdated dependencies using dependency update metrics face two key challenges.
@@
1. Metric Limitations. The calculations used by existing dependency update metrics often do not handle floating version constraints or are not designed to localize vulnerabilities (coxMeasuringDependencyFreshness2015, ; gonzalez-barahona_technical_2017, ). Existing metrics also often calculate an update time for each of a dependency and for each version of a package, making it difficult to understand that package’s update practices as a whole. Additionally, existing metrics do not weight the recent dependency update practice which makes the existing metrics less actionable for developers. Finally, there are no existing publicly available dependency update metrics specifically for vulnerable dependencies. While some industry reports use MTTR in this way, their specific calculations are not available.
2. Insufficient Externally Reported Vulnerability Data. A small fraction of packages in software ecosystems have reported vulnerabilities. For example, 1363 npm, 694 PyPI, and 383 Cargo (total 2.4K) packages have externally-reported vulnerabilities as of 2024-09-12. Practitioners can only compute MTTR for the dependents of those 2.4K packages. 0.49% npm, 0.83% PyPI, and 0.05% Cargo packages have directly depended on at least one of these 2.4k vulnerable packages in their lifetime. In addition, the lack of externally reported vulnerability data is also discussed by related research (alhazmi2007measuring, ; zahan_software_2023, ).
The goal of this study is to aid practitioners in understanding how quickly packages update their dependencies through an empirical study using two novel dependency update metrics. With this goal, we conduct our study with four research questions.
RQ1::
How do we measure the MTTU and MTTR of a package including its dependencies such that the measure is responsive to more recent update practices and therefore actionable to the development team?
We first propose novel algorithms for computing Mean-Time-To-Update (MTTU) and Mean-Time-To-Remediate (MTTR), which we denote MTTUdep and MTTRdep, respectively, to overcome challenges with existing dependency update metric calculations.
After defining the two novel metrics, we performed an empirical analysis on npm, PyPI, and Cargo ecosystems. We collect package version release information and dependency relations of packages of $163,207$ packages from the three ecosystems and compute MTTUdep and MTTRdep. With the computed metrics, we answer the second research question.
RQ2::
How do packages in npm, PyPI, and Cargo differ in MTTU and MTTR?
We analyze the distributions of MTTUdep and MTTRdep using violin plots to explore the differences in the ecosystems.
In all three ecosystems, MTTRdep cannot be computed for 99.51% npm, 99.17% PyPI, and 99.95% Cargo packages since these packages have not depended upon any vulnerable direct dependencies in their lifetime. While dependency update metrics are useful, this lack of externally reported vulnerability data limits their application. To provide dependency update measurements for those 99%+ of packages without externally-reported vulnerabilities, we analyzed whether MTTUdep would provide a practical estimation/ proxy. This analysis is inspired by Dhrymes and Guerard (dhrymes1978introductory, ), who suggested that when a variable is unobservable, a proxy variable, which is a variable that can be used as a substitute for the missing one, can be used. This leads us to our third research question.
RQ3::
Can MTTU be used as a proxy for MTTR?
We perform a proxy analysis, consisting of a set of statistical tests from literature, on MTTUdep and MTTRdep to explore if MTTUdep can serve as a proxy for MTTRdep.
Finally, we would like to understand which package characteristics (e.g., contributors count, version count) have more influence on MTTUdep and MTTRdep, which leads to our last research questions.
RQ4::
How do package characteristics influence MTTU and MTTR?
We use correlation tests to understand the association between nine package characteristics and MTTUdep/ MTTRdep. We further conduct regression analysis to quantify which package characteristics matter more in influencing MTTUdep and MTTRdep values.
@@
Contributions. In summary, this paper contributes (1) a detailed algorithm and process for quantifying the dependency update practice of a package using our novel metrics; (2) statistical hypothesis testing on using MTTU as a proxy for MTTR, when externally reported vulnerability data is not available; (3) a large-scale analysis of the dependency update metrics in npm, PyPI, and Cargo packages; and (4) correlation and regression analysis illustrating which package characteristics impact the likelihood of higher MTTU and MTTR.
@@
We provide our replication package in Zenodo (zenodo-artifact, ), currently restricted for reviewers only. Upon acceptance of the paper, we will make it public.
2. Background And Related Work
@@
In this section, we provide a brief overview of the existing update metric and discuss related work.
@@
Further research by Zerouali et al. (zeroualiFormalFrameworkMeasuring2019, ) propose a formal framework for measuring technical lag in software ecosystems. They analyzed 4M releases of 500K npm packages, considering the evolution of technical lag over time. They found that technical lag induced by direct dependencies in npm packages increases over time due to missed updates, including major releases. Stringer et al. (stringerTechnicalLagDependencies2020, ) study the technical lag of dependencies in a large-scale cross-ecosystem fashion containing packages from 14 package managers. They found that pinned dependencies are the main reason behind technical lag. Zerouali et al. (zeroualiMultidimensionalAnalysisTechnical2021, ) expand the idea of ‘technical lag’ into multiple dimensions (package lag, time lag, version lag, vulnerability lag, and bug lag) and study the technical lag in 140K Docker images. They found that the median time lag of community Docker images is over a year. Although previous studies have explored technical lag in terms of general dependency updates, understanding vulnerable dependency update practice was not their goal. In contrast, we study the dependency update metric and the vulnerable dependency update metric and investigate their relationship and other characteristics.
@@
Pashchenko et al. (pashchenkoVulnerableOpenSource2018, ) studied the most used Java dependencies in SAP software and found that only updating the dependencies’ version can remove 81% vulnerable dependencies. Kula et al. (kulaDevelopersUpdateTheir2018, ) studied the update behavior of developers w.r.t. security advisories and found that developers do not update their vulnerable dependencies regularly. Kumar et al. (kumar_comprehensive_2024, ) conducted a study to understand how widespread vulnerabilities are and how quickly they are being fixed. They found that for most programming languages, a critical vulnerability persists on average for over a year before being fixed. Studies in outdated dependencies and vulnerable dependencies are focused on either all updates or only security updates, but not both. In contrast, we focus on both outdated and vulnerable dependencies and explore their relationship using our proposed metrics.
MTTU and MTTR. The metrics MTTU and MTTR have been used in the software reliability and maintenance domain for a long time (gokhaleTimeStructureBased1999, ; calvoApplyingGoalQuestion2023, ; johnsonSurveySoftwareTools1988, ; morrisonMappingFieldSoftware2018, ). Researchers have also studied different security metrics (time to close bug/vulnerability, window of exposure, vulnerability count) (davidsonGoodBadUgly2009, ; beresUsingSecurityMetrics2009, ) of various categories (time metric, vulnerability metric) in the software security domain, but these metrics are focused on vulnerabilities in packages [^1] Faults and bugs are considered in MTTR instead of vulnerability in Reliability domain research.. In our study, we focus on measuring dependency update metrics for packages having vulnerable dependencies, not the vulnerable package itself.
@@ 3. Challenges Applying Existing Update Metrics
$0.9.2:\operatorname{2022-03-22}$ $^{\wedge}3.4.0$ $3.4.0:\operatorname{2022-03-18}$ $3.4.0:\operatorname{2022-03-18}$ 0
@@ 3.1. Technical Lag Using tLag
Time lag (tLagtLag(zerouali_empirical_2018, ; decan_evolution_2018, ) is a way to measure technical lag to assess the outdatedness of a dependency in terms of time. Conceptually, $tLag$ measures the time difference between the release time of the latest version of a dependency and the release time of the version of the dependency used by a package, at the time the package was released.
Let pkgpkg be a package, and $pkg_{v}$ indicate a specific version $v$ of pkgpkg. Let depdep be a direct dependency of pkgpkg. When specifying depdep as a dependency, $pkg_{v}$ specifies a dependency constraint $pkg_{v}.dep.\bar{c}$. Package managers use dependency constraints to resolve the highest version that satisfies the constraint. For example, a dependency constraint $pkg_{v}.dep.\bar{c}$ of $^{\wedge}1.2.3$” is satisfied for the range [¿=1.2.3, ¡ 2.0.0]. We use the function $lastAllowed(dep,\bar{c},t)$ to denote the version resolution of depdep for constraint $\bar{c}$ at time $t$. We further use the function $latest(dep,t)$ to denote the latest version of depdep at time $t$ and $time(\cdot)$ to denote the time a given version is released. Hence, the $tLag$ of $pkg_{v}$ for depdep is formally defined in Equation 1.
@@
Table 1 shows an example computation of technical lag for codemod-cli’s dependency on simple-git. The first two columns of Table 1 indicate the version with the date of different releases of package codemod-cli and the version constraint specified for dependency simple-git. The “lastAllowed(simple-git)” column represents the resolved version of dependency simple-git with specified constraint at the time of the release of codemod-cli. For computing the technical lag, we need to know the latest available version of the dependency simple-git at the time of the release of each version of codemod-cli according to Eq. 1, which is represented in column “latest(simple-git)”. The tLag column is the computed technical lag. We subtract the release date of lastAllowed(simple-git) from the release date of latest(simple-git) to obtain the technical lag for each release of codemod-cli. The computed tLag of five versions of codemod-cli were 707, 78, 78, 100, and 0, consecutively, in Table 1 and Fig. 1.
@@ 3.2. Design Gaps
Package-Level Metric. A dependency update metric should combine all data points into one single value per package to make it easier for developers to understand and compare metrics. However, tLag computes the technical lag for each version of the package for a dependency as shown in Table 1. For developers, understanding what to do with all of these tLag values for a package with multiple releases and multiple dependencies is hard. Having an aggregation would improve the utility of dependency update metrics.
@@
Localizing Vulnerabilities. A security-oriented dependency update metric should incorporate the update characteristics of vulnerable dependencies. tLag includes bug fixes, feature updates, and security fixes as a whole. Technically, tLag’s measurement can be modified to only consider the dependency’s security fixes adopted by the package in its formal framework. However, no previous work measured tLag for measuring vulnerable dependency update practices.
4
Weighting Lifetime And Recent Practices. A dependency update metric should reflect more heavily the most recent dependency updates made by the package developers. However, tLag does not weight recency into its measurement. Recent update practice provides the most relevant information for ongoing maintenance of the package. For example, the development team’s update practice from 10 years ago should not be weighted as heavily as their current update practice. In addition, recency makes a dependency update metric more actionable since a metric with recency can be used to compare two packages with different lifespans (e.g., two packages with the same functionality, one with a shorter lifespan and the other with a longer lifespan). Actionability is a desired property for good software metrics (meneely_validating_metrics_2013, ).
4. RQRQ1:: Novel Dependency Update Metrics